OpenStudy (anonymous):

Help on finding the critical points of the following function with its first derivative f'(x)=2e^x(x^2+25)(x^2-16)(x-3)^8 --------------------------- ln(x^2+100)

OpenStudy (anonymous):

*

OpenStudy (anonymous):

Taking the derivative of this thing is a pain. I'd use mathematica, personally.

OpenStudy (anonymous):

This is already the first deriviative of a function

OpenStudy (anonymous):

Im having trouble on ow to solve for X

OpenStudy (anonymous):

Oh, I see.

OpenStudy (anonymous):

Because that is how you find ciritcl points right?

OpenStudy (anonymous):

The critical points are any x values that make the first derivative 0.

OpenStudy (anonymous):

So if you set the derivative equal to 0, then solve for x, that should do it for ya.

OpenStudy (anonymous):

It cant be a a whole number because the denominator is a log. It would have to be E or something along the lines of that

OpenStudy (anonymous):

Oh, and it's also a critical point if the denominator is 0.

OpenStudy (anonymous):

Not true. If the numerator is 0, then the whole thing will be 0.

OpenStudy (anonymous):

0/anything = 0

OpenStudy (anonymous):

Except for 0/0 which is undefined.

OpenStudy (anonymous):

Ok. So I started off by multiplying by the denominator for both sides. Is that correct?

OpenStudy (anonymous):

Sure. What equation did that yield you?

OpenStudy (anonymous):

set \[2e^x(x^2+25)(x^2-16)(x-3)^8=0\]

OpenStudy (anonymous):

as a fraction is only zero when the numerator is then note that \(e^x\) is never zero neither is \(x^2+25\) so it becomes much easier

OpenStudy (anonymous):

\[\ln (x^2+100)= 2e^x(x^2+25)(x^2-16)(x-3)^8\]

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Incorrect.

OpenStudy (anonymous):

woops sorry. The left half hould be 0 because it is multiplied by 0

OpenStudy (anonymous):

you are not trying to find where it is one, only where it is zero

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

forget the denominator, it is never zero, just concentrate on the numerator

OpenStudy (anonymous):

Ok, I first divided by 2 on both sides I got \[e^x(x^2+25)(x^2-16)(x-3)^8\] Then i took the 8th root I got \[e^x(x^2+25)(x^2-16)(x-3)\]

OpenStudy (anonymous):

That actually doesn't work, algebraically speaking. You would have to take the 8th root of the whole side, not just the one term. Instead, consider this fact. That side is just 4 things multiplied together. If any one of those 4 things is 0, then the whole side will be 0.

OpenStudy (anonymous):

Now, recognizing that, just look at each of the 4 things and ask yourself when that thing will be 0.

OpenStudy (anonymous):

Oh since (x^2-16 cant be 4, a critical point is 4

OpenStudy (anonymous):

and 3

OpenStudy (anonymous):

Well, I'm not sure what you mean by saying it "can't be 4" but if x=4, then (x^2-16) would be 0, so the whole thing would be 0.

OpenStudy (anonymous):

Right. 3 is another critical point.

OpenStudy (anonymous):

So 0/3/4 are all critical points. Wouldnt -4 also be one as well?

OpenStudy (anonymous):

0 is actually not a critical point. I'm not sure why you just said it was.

OpenStudy (anonymous):

But -4 is one.

OpenStudy (anonymous):

Ok so how would I write this out if I am supposed t show ym steps? I just ignore the denominator and state that since if the numerator is 0, the whole answer will be 0?

OpenStudy (anonymous):

Essentially. It's not hard to show that. 0/anything = 0

OpenStudy (anonymous):

Or, if you want to do it algebraically, you can just do what you actually did on your own. Set the whole thing equal to 0, then multiply both sides by the denominator.

OpenStudy (anonymous):

My professor is a real stickler when it comes to work. But since that I stated that 0/anything is always 0, he cant be a jerk about it.

OpenStudy (anonymous):

If he's a stickler, just do what you did on your own. That's not anything he can complain about. \(\Large \frac{2e^x(x^+25)(x^2-16)(x-3)^8}{ln(x^2+100)}=0\) Multiply both sides by \(ln(x^2+100)\)

OpenStudy (anonymous):

gives \(2e^x(x^2+25)(x^2-16)(x-3)^8 =0\)

OpenStudy (anonymous):

Then divide by 2 which gives \[e^x(x^2+25)(x^2-16)(x-3)^8\]

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

Satellite73, could you help me solve this equation algebraically?